# Holographic inflation and holographic dark energy from entropy of the anti-de Sitter black hole

### v1

@@ III.1 Evolution and stability
Thus, according to the above results, since both $q$ and $\omega_{de}$ approach $-1$, the current acceleration can be realized, and NHDEH can behave as the cosmological constant. The universe will eventually evolve into an epoch described by the standard $\Lambda$CDM model, and the whole evolution of the universe can be described by this model, which is depicted in the right panel of Fig. 6.
III.2 Dynamical analysis
In the previous subsection, we analyzed the evolution of the universe in NHDEH and found it to be stable against perturbations. In this subsection, we will analyze the dynamical behavior of NHDEH. In order to investigate the complete asymptotic behavior of NHDEH, we will adopt dynamical system analysis Bahamonde2018 ; Wu2010 ; Dutta2017 ; Huang2019a ; Huang2021 . The critical points can be obtained by solving the equations of the autonomous system
@@ III.2 Dynamical analysis
Linearizing the equations of the autonomous system  (40) and  (41), we obtain the corresponding first order differential equations. The coefficient matrix of these equations determines the stability of the critical points, which are presented in Table 1. For a critical point, if all eigenvalues are negative, it is a stable point representing an attractor; if all eigenvalues are positive, the corresponding point is unstable; if the eigenvalues have opposite signs, the critical point is a saddle point. From Table 1, we can see that point $P_{1}$ is unstable, $P_{2}$ is a saddle point, and $P_{3}$ is stable. According to the stability of these critical points, we can see that the universe evolves from the radiation dominated epoch $(P_{1})$ into the pressureless matter dominated epoch $(P_{2})$, and ultimately enters the NHDEH dominated late-time acceleration epoch $(P_{3})$. The behavior of attractor $P_{3}$ and the evolutionary trajectories are shown in Fig. 7, where the left panel is plotted for $b^{2}=0.01$ and the right one is for $b^{2}=0.1$. Compared to the right panel of Fig. 7, the left one is more suitable for describing the entire evolutionary history of the universe. Therefore, a smaller $b^{2}$ is required to describe the entire evolutionary history of the universe in NHDEH, and the universe will eventually evolve into an epoch characterized by the cosmological constant $\Lambda$.
@@ III.3 Statefinder analysis
The statefinder parameters $r$ and $s$, which are geometrical diagnostics and depend only on $a$, are defined as Sahni2003
@@
Differentiating Eq. (38), the statefinder parameters $r$ and $s$ can be expressed in terms of $\Omega_{m}^{\prime}$ and $\Omega_{de}^{\prime}$. Thus, we can obtain the evolution curves of the universe in the $r-s$ parameter space by numerically solving Eqs. (40) and (41). In the left panel of Fig. (8), we plot an example for the evolution curves of the statefinder diagnostic pair $\{r,s\}$.
In this panel, the blue line represents the $\Lambda$CDM model, while the red, green, and purple dashed lines denote the evolutionary curves for $b^{2}=0.01$, $0.1$, and $0.5$; the blue dot marks the $\Lambda$CDM fixed point $(0,1)$, and the red, green, and purple dots represent the current values for $b^{2}=0.01$, $0.1$, and $0.5$. The results from the left panel of Fig. (8) indicate that we cannot distinguish NHDEH from the standard $\Lambda$CDM model using the statefinder diagnostic pairs $\{r,s\}$, as all the evolutionary curves and points overlap. Then, we depict another statefinder diagnostic pairs $\{r,q\}$ in the right panel of Fig. (8), in which the blue dot denotes the de Sitter expansion fixed point $(-1,1)$ in the future and the orange dot represents the standard cold dark matter fixed point $(0.5,1)$. This panel shows that all the curves overlap, starting from the standard cold dark matter fixed point $(0.5,1)$ and then evolving into the de Sitter expansion fixed point $(-1,1)$. Thus, using the statefinder diagnostic pairs $\{r,s\}$ and $\{r,q\}$, we cannot distinguish NHDEH from the standard $\Lambda$CDM model.
@@
III.4 Hubble diagram
By considering the future event horizon as an IR cutoff, it is found that there exists a turning point in the Hubble diagram $H(z)$ in the holographic dark energy model, where the evolutionary curves of the Hubble parameter with the cosmological paradigm Colgain2021 . Subsequently, it is found that this turning point can be avoided in the Barrow holographic dark energy model under some specific circumstances Huang2021 . Lastly, we will discuss whether this turning point exists in NHDEH in this subsection.
@@ IV Conclusion
Subsequently, we analyze the evolution of the universe in this model and then apply dynamical analysis techniques to it. Our study indicates that the late-time acceleration of the universe can be achieved as $q$ and $\omega_{de}$ approach $-1$. This model is stable against perturbations, as the squared sound speed $v^{2}_{s}$ is positive. It can describe the entire evolutionary history of the universe when $b^{2}$ takes small values, with radiation, pressureless matter, and NHDEH successively dominate the evolution of the universe. Due to the presence of an attractor corresponding to the epoch described by the $\Lambda$CDM model, the universe will eventually evolve into an epoch characterized by the cosmological constant $\Lambda$. To distinguish this model from the standard $\Lambda$CDM model, we apply the statefinder analysis method. When using the statefinder diagnostic pairs $\{r,s\}$ and $\{r,q\}$ to discriminate this model from the standard $\Lambda$CDM model, we find that all the evolutionary curves and points overlap in the $r-s$ plane, and all the evolutionary curves overlap in the $r-q$ plane. Therefore, it is difficult to distinguish this model from the standard $\Lambda$CDM model using the statefinder diagnostic pairs. Then, we discuss the turning point of the Hubble diagram in this model. We find that the evolutionary curves of the Hubble parameter $H$ overlap with the curve of the standard $\Lambda$CDM model for sufficiently small values of $b^{2}$, and the turning point disappears when $b^{2}$ is small. In particular, for $b^{2}=0.01$, the turning point does not exist. These results indicate that when $b^{2}$ takes a small value, NHDEH cannot be distinguished from the standard $\Lambda$CDM model and can serve as an alternative model to the standard $\Lambda$CDM model.

### v2

@@ IV.1 Evolution and stability
Thus, according to the above results, since both $q$ and $\omega_{de}$ approach $-1$, the current acceleration can be realized, and NHDEH can behave as the cosmological constant. The universe will eventually evolve into an epoch described by the standard $\Lambda$CDM model, and the whole evolution of the universe can be described by this model, which is depicted in the right panel of Fig. (8).
IV.2 Dynamical analysis
In the previous subsection, we analyzed the evolution of the universe in NHDEH and found it to be stable against perturbations. In this subsection, we will analyze the dynamical behavior of NHDEH. In order to investigate the complete asymptotic behavior of NHDEH, we will adopt dynamical system analysis Bahamonde2018 ; Wu2010 ; Dutta2017 ; Huang2019a ; Huang2021 ; Wu2007 ; Wu2008 . The critical points can be obtained by solving the equations of the autonomous system
@@ IV.2 Dynamical analysis
Linearizing the equations of the autonomous system  (58) and  (59), we obtain the corresponding first order differential equations. The coefficient matrix of these equations determines the stability of the critical points, which are presented in Table 1. For a critical point, if all eigenvalues are negative, it is a stable point representing an attractor; if all eigenvalues are positive, the corresponding point is unstable; if the eigenvalues have opposite signs, the critical point is a saddle point. From Table 1, we can see that point $P_{1}$ is unstable, $P_{2}$ is a saddle point, and $P_{3}$ is stable. According to the stability of these critical points, we can see that the universe evolves from the radiation dominated epoch $(P_{1})$ into the pressureless matter dominated epoch $(P_{2})$, and ultimately enters the NHDEH dominated late-time acceleration epoch $(P_{3})$. The behavior of attractor $P_{3}$ and the evolutionary trajectories are shown in Fig. (9), where the left panel is plotted for $b^{2}=0.01$ and the right one is for $b^{2}=0.1$. Compared to the right panel of Fig. (9), the left one is more suitable for describing the entire evolutionary history of the universe. Therefore, a smaller $b^{2}$ is required to describe the entire evolutionary history of the universe in NHDEH, and the universe will eventually evolve into an epoch characterized by the cosmological constant $\Lambda$.
@@ IV.3 Statefinder analysis
The statefinder parameters $r$ and $s$, which are geometrical diagnostics and depend only on $a$, are defined as Sahni2003 ; Wu2005
@@
Differentiating Eq. (56), the statefinder parameters $r$ and $s$ can be expressed in terms of $\Omega_{m}^{\prime}$ and $\Omega_{de}^{\prime}$. Thus, we can obtain the evolution curves of the universe in the $r-s$ parameter space by numerically solving Eqs. (58) and (59). In the left panel of Fig. (10), we plot an example for the evolution curves of the statefinder diagnostic pair $\{r,s\}$. In this panel, the blue line represents the $\Lambda$CDM model, while the red, green, and purple dashed lines denote the evolutionary curves for $b^{2}=0.01$, $0.1$, and $0.5$; the blue dot marks the $\Lambda$CDM fixed point $(0,1)$, and the red, green, and purple dots represent the current values for $b^{2}=0.01$, $0.1$, and $0.5$. The results from the left panel of Fig. (10) indicate that we cannot distinguish NHDEH from the standard $\Lambda$CDM model using the statefinder diagnostic pairs $\{r,s\}$, as all the evolutionary curves and points overlap. Then, we depict another statefinder diagnostic pairs $\{r,q\}$ in the right panel of Fig. (10), in which the blue dot denotes the de Sitter expansion fixed point $(-1,1)$ in the future and the orange dot represents the standard cold dark matter fixed point $(0.5,1)$. This panel shows that all the curves overlap, starting from the standard cold dark matter fixed point $(0.5,1)$ and then evolving into the de Sitter expansion fixed point $(-1,1)$. Thus, using the statefinder diagnostic pairs $\{r,s\}$ and $\{r,q\}$, we cannot distinguish NHDEH from the standard $\Lambda$CDM model.
@@
IV.4 Hubble diagram
By considering the future event horizon as an IR cutoff, it is found that there exists a turning point in the Hubble diagram $H(z)$ in the holographic dark energy model Colgain2021 . Subsequently, it is found that this turning point can be avoided in the Barrow holographic dark energy model under some specific circumstances Huang2021 . In this subsection, we will discuss whether this turning point exists in NHDEH.
@@ V Conclusion
Subsequently, we investigate reheating in this model by calculating the reheating temperature and the number of reheating e-folds. Our results demonstrate that instantaneous reheating is excluded by the Planck results, as the corresponding spectral index $n_{s}$ exceeds the $2\sigma$ bound on $n_{s}$ from Planck data. For the cases $n=\frac{1}{3}$ and $n=\frac{1}{2}$ with the observed value $b^{2}=0.01$, the allowed ranges for the reheating temperature $T_{reh}$ and the number of reheating e-folds $N_{reh}$ are $10^{-2}GeV<T_{reh}<10^{5}GeV$ and $429<N_{reh}<$ 42, respectively. These results are consistent with both theoretical predictions and observational constraints.
Finally, we analyze the evolution of the universe in this model and then apply dynamical analysis techniques to it. Our study indicates that the late-time acceleration of the universe can be achieved as $q$ and $\omega_{de}$ approach $-1$. This model is stable against perturbations, as the squared sound speed $v^{2}_{s}$ is positive. It can describe the entire evolutionary history of the universe when $b^{2}$ takes small values, with radiation, pressureless matter, and NHDEH successively dominate the evolution of the universe. Due to the presence of an attractor corresponding to the epoch described by the $\Lambda$CDM model, the universe will eventually evolve into an epoch characterized by the cosmological constant $\Lambda$. To distinguish this model from the standard $\Lambda$CDM model, we apply the statefinder analysis method. When using the statefinder diagnostic pairs $\{r,s\}$ and $\{r,q\}$ to discriminate this model from the standard $\Lambda$CDM model, we find that all the evolutionary curves and points overlap in the $r-s$ plane, and all the evolutionary curves overlap in the $r-q$ plane. Therefore, it is difficult to distinguish this model from the standard $\Lambda$CDM model using the statefinder diagnostic pairs. Then, we discuss the turning point of the Hubble diagram in this model. We find that the evolutionary curves of the Hubble parameter $H$ overlap with the curve of the standard $\Lambda$CDM model for sufficiently small values of $b^{2}$, and the turning point disappears when $b^{2}$ is small. In particular, for $b^{2}=0.01$, the turning point does not exist. These results indicate that when $b^{2}$ takes a small value, NHDEH cannot be distinguished from the standard $\Lambda$CDM model and can serve as an alternative model to the standard $\Lambda$CDM model.